# Efficiently Editing Mixture-of-Experts Models with Compressed Experts

### v1

@@ 1 Introduction
Figure 1: The inference time of MoE models grows linearly with increasing number of activated experts. The plot is generated using OLMoE architecture. Details on the inference speed measurement methodology are provided in Section 3.1.
In practice, MoE models often require multiple activated experts to achieve desirable performance. For instance, Mixtral (Jiang et al., 2024) and Phi-MoE (Abdin et al., 2024) activate two experts, Qwen-MoE (Yang et al., 2024) activates four and OLMoE (Muennighoff et al., 2024) activates eight. The efficiency of MoE models largely depends on the number of activated experts. Specifically, the inference cost of MoE models increases linearly with this number (as shown in Figure 1). Fewer activated experts reduce training and deployment costs as a smaller portion of the network is used. Conversely, activating more experts increases utilization of model capacity, often leading to superior performance. Thus, the selection of the number of activated experts poses a fundamental trade-off between efficiency and performance.
@@ 2.1 MoE Layers
where $\alpha_{i}=\text{top-}k(\text{Softmax}(R(h)))_{i}$. In this formulation, the top-$k$ mechanism ensures sparse activation, where only the $k$ experts with the highest routing weights contribute to the output, while the remaining $n-k$ experts are effectively disabled ($\alpha_{i}=0$). Typically, only a small fraction of experts is utilized during each forward pass, i.e., $k\ll n$, enabling substantial computational savings. For instance, Phi-MoE (Abdin et al., 2024) activates 2 out of 16 experts, and OLMoE (Muennighoff et al., 2024) activates 8 out of 64 experts, etc. This sparse activation significantly reduces computational costs compared to dense models, making MoE layers efficient and scalable.
@@ 2.2 Compressed Experts
Conventionally, for each input token, $k$ out of the $n$ experts are activated. We can categorize these activated experts into $k_{m}$ main experts and $k_{a}=k-k_{m}$ auxiliary experts, based on their ordered routing weights. In many scenarios, instead of fully activating all $k$ experts, only forwarding the main experts can lead to a better performance-cost trade-off. However, we do not want to completely discard the information in auxiliary experts. Compressed experts provide a solution to leverage their compressed counterparts to reduce computational cost while retaining their contributions. We introduce it with the following three steps (illustrated in Figure 3):
  1. 1.
@@
  3. 3.
    Hidden states augmentation: The aggregated compressed expert $\theta$ is incorporated into the hidden state $h$ through an element-wise product. This operation has proven highly effective by PEFT methods such as (IA)3 (Liu et al., 2022), which applies a similar transformation to attention activations. Then, the augmented hidden state is passed only through the main experts, eliminating the need for forward computation through the $k_{a}$ auxiliary experts, thereby reducing inference cost.
@@ 2.3 Expert Reduction
While compressed experts are flexible and can be applied to any number of main and auxiliary experts, we investigate the conditions under which they are most effective. Specifically, we use OLMoE, which has 8 active experts ($k=8$), as a case study. We evaluate different configurations by varying the number of main experts ($k_{m}\in\{1,2,4,8\}$), and replace the remaining $k_{a}=k-k_{m}$ auxiliary experts with compressed experts (abbreviated as CE). To assess the impact of compressed experts in downstream adaptation, we conduct experiments using supervised finetuning (SFT) on pretrained OLMoE with mathematical data, followed by evaluation on GSM8k with 0-shot CoT (details provided in Section 3.1).
@@
As shown in Table 1, compressed experts consistently improve performance across reduced configurations, demonstrating their ability to recover performance lost due to expert reduction. However, their effectiveness is limited when more than half of the experts are treated as auxiliary.
@@ 2.4 Reduction in Activate Parameters
Replacing half of the activated experts with compressed experts significantly reduces the number of active parameters. Taking Phi-MoE as an example, each MoE layer consists of three weight matrices of size $4096\times 6400$. The total non-MoE parameters are approximately 2.4B. In the original configuration with 2 activated experts across 32 MoE layers, the total number of MoE parameters is $3\times 4096\times 6400\times 32\times 2\approx 5.03B.$ With compressed experts, only 1 expert is activated per layer, and the compressed experts add minimal parameters (underlined below): $3\times(4096\times 6400\times 1+\underline{2\times 4096+6400})\times 32\approx 2.52B$. In total, in one forward pass, the number of active parameters for the original configuration is 7.45B, while the one with compressed experts is only 4.93B, resulting in a saving of 33.8%. We provide a similar calculation for OLMoE in Appendix A, which results in a 31.4% reduction in active parameters.
Figure 4: The performance of Phi-MoE versus the inference latency, each point representing a different expert configuration. The Top-1 w/ CE configuration performs closely to the Top-2 configuration while achieving low inference latency close to Top-1.
@@ 2.5 Performance-Latency Trade-off
To analyze the effectiveness of compressed experts, we compare the performance and inference latency of three configurations: top-2, top-1 with compressed experts (abbreviated as CE), and top-1, as shown in Figure 4. The performance is evaluated on the Phi-MoE finetuned on mathematical reasoning data, with detailed training and evaluation procedures described in Section 3.1.
@@
Introducing compressed experts in the top-1 w/ CE configuration effectively addresses this trade-off. With a latency of 4.35 seconds, the compressed expert configuration adds only a minimal computational overhead compared to top-1. However, it achieves a substantial performance improvement over top-1, closing much of the gap between top-1 and top-2 by reaching a performance of 57.4%. This demonstrates the ability of compressed experts to augment the utility of the top-1 expert with minimal additional cost, making it a practical and efficient model editing approach between performance and latency.
@@ 3 Experiments
In Section 3.1, we detail the experimental setup, including model configurations, datasets and training specifics. In Section 3.2, we first present the performance of compressed experts on evaluation benchmarks, then show their the inference cost savings, demonstrating that compressed experts achieves an effective balance between performance and efficiency. Finally, in Section 3.3, we provide a detailed analysis on the task-specific performance of compressed experts.
@@ 3.1 Setup
Models
We evaluate our approach on two representative MoE models: Phi-MoE (Abdin et al., 2024) and OLMoE (Muennighoff et al., 2024). These models vary in size and the number of activated experts, as summarized in Table 2.
Training details
Task ()($\rightarrow$)( → )
Metric ()($\rightarrow$)( → )
@@
Task ()($\rightarrow$)( → )
Metric ()($\rightarrow$)( → )
@@
Task ()($\rightarrow$)( → )
Metric ()($\rightarrow$)( → )
@@
Task ()($\rightarrow$)( → )
Metric ()($\rightarrow$)( → )
@@
Training datasets
To evaluate compressed experts in various domains and applications, we conduct experiments on three tasks with different scopes. For specialized tasks including mathematics and coding, we finetune models on MathInstruct (Yue et al., 2024) and Magicoder (Wei et al., 2023). For general capabilities, we finetune models on TÜLU 3 (Lambert et al., 2024), a general-purpose instruction-following dataset targeting skills including reasoning, math, coding and safety.
Evaluation datasets
For evaluation, we benchmark models on both specialized and general tasks. For mathematical reasoning and coding, we evaluate on GSM8K (Cobbe et al., 2021) and HumanEval (Chen et al., 2021) respectively. To assess general-purpose capabilities, we further include IFEval (Zhou et al., 2023), TruthfulQA (Lin et al., 2021), BBH (Suzgun et al., 2023). Note that for HumanEval, we use Pass@1 for Phi-MoE and Pass@10 for OLMoE respectively. Since OLMoE is a smaller model with inherently lower capacity for complex coding tasks, Pass@10 provides finer-grained resolution to better capture performance differences across various configurations. The detailed evaluation metric for each task are presented in the tables shown below.
Baselines
Following the methodology outlined in Section 2.5, we replace half of the activated experts with compressed experts and compare their performance against two baselines: the pretrained model and the configuration with half of the activated experts. Additionally, we report the performance of the full-expert configuration as a performance oracle for reference.
Inference latency evaluation
To evaluate the reduction of inference time with compressed experts, we follow the practice in the latency benchmark presented in Kwon et al. (2023). We create dummy prompts of batch size 8 and sequence length 32 with randomly generated token IDs. Then, we pass the dummy prompts to the model and let it generate completions. The model first goes through 10 warmup iterations, and the reported latency is averaged over 30 iterations of completions.
@@ 3.2 Main Results
We present the performance of compressed experts on general tasks and specialized tasks respectively. Table 3 and Table 4 present the performance on a diverse set of general evaluation tasks, along with the inference latency for each configuration. Table 5 and Table 6 present the performance of Phi-MoE and OLMoE on specialized tasks, including mathematic and coding. Beyond raw performance, we additionally report normalized performance: $\text{perf}_{\text{CE}}/\text{perf}_{\text{full}}$. It quantifies how well compressed experts approximate the performance of the full-expert configuration, providing a direct measure of performance retention after expert reduction.
Performance comparison
The results demonstrate that compressed experts significantly narrow the performance gap with fully activated expert configurations. Across all models and tasks, compressed experts consistently recover more than 90% of the full expert performance, showcasing their ability to retain key information from auxiliary experts. Additionally, the incorporation of compressed experts consistently outperform the halved baseline. This shows that information contained in auxiliary experts still has a meaningful contribution to the model performance. This allows MoE models to maintain strong task performance with fewer active parameters. Similar to the findings in Huang et al. (2024), we observe that the optimal number of activated experts varies based on task complexity and broadness, and compressed experts exhibit task-dependent efficacy, which we detail in Section 3.3.
Inference latency
Augmenting halved configurations with compressed experts (CE) incurs only marginal overhead. For Phi-MoE, top-1 with CE increases latency by 8.5% over top-1, while OLMoE’s top-4 with CE adds 9.7% latency compared to vanilla top-4. This overhead arises from lightweight operations such as element-wise multiplications, which is orders of magnitude cheaper than executing a full expert. Critically, CE-augmented configurations remain far more efficient than their original counterparts: Phi-MoE top-1 with CE is 22% faster than top-2, and OLMoE top-4 with CE is 18.4% faster than top-8. Combined with their strong empirical performance, we comprehensively validate that the compressed experts effectively balance efficiency and performance.
@@ 3.3 Empirical Analysis
We provide a more detailed analysis on per-task performance in the general evaluation setting (Tables 3 and 4). Unlike specialized evaluation with a focus on specific domains, this setting covers factual correctness, logical reasoning, and precise instruction-following abilities. The diversity of tasks makes it challenging for MoE models to optimize for every task simultaneously, which in turn affects how well compressed experts perform. We observe that not all tasks benefit equally from compressed experts. Here, we mainly base our analysis on the performance improvement from the halved baseline.
@@
Overall, compressed experts provide a highly effective balance between efficiency and performance, consistently recovering over 90% of the full expert performance across both models. The performance recovery could vary across different task types, with structured reasoning tasks benefiting more than knowledge-intensive or multi-step reasoning tasks.
@@ 4 Related Works
Mixture-of-Experts (MoE) models have emerged as a powerful approach to scaling up model capacity while maintaining computational efficiency. By activating only a subset of experts during inference, MoE models significantly reduce computational overhead compared to dense models. This advantage has led to a surge in open-source MoE implementations, each varying in the number of activated experts per layer. For instance, Mixtral (Jiang et al., 2024), JetMoE (Shen et al., 2024), OpenMoE (Xue et al., 2024) and Phi-MoE (Abdin et al., 2024) activate two experts, Qwen-MoE (Yang et al., 2024) activates four, DeepSeekMoE (Dai et al., 2024) activates six and OLMoE (Muennighoff et al., 2024) activates eight. However, recent work by Huang et al. (2024) demonstrates that fully utilizing all activated experts can introduce redundancy for certain tasks, as not all tasks require the full capacity of the top-$k$ experts.
Inspired by this insight, we propose compressed experts, designed to compress auxiliary experts during model editing. Compressed experts are model-agnostic and can be seamlessly integrated into existing MoE architectures, including those mentioned above. By reducing the number of activated experts while maintaining performance, our approach enhances computational efficiency without compromising model quality.
Soft merging of experts in MoE
@@
Inspired by the surge of model merging techniques (Ilharco et al., 2023; Wortsman et al., 2022; Yadav et al., 2023; Yang et al., 2023; He et al., 2025), recent works propose to merge activated experts for more precise gradient computation and faster inference. However, these approaches suffer from a common limitation: the merging operations introduce significant computational overhead, making them impractical for token-based routing. Specifically, those approaches require merging experts on-the-fly based on the per-token routing weights. For instance, SMEAR (Muqeeth et al., 2024) and Lory (Zhong et al., 2024) merge experts via averaging the expert parameters based on the routing weights. While effective in sequence-based or semantic-based routing scenarios, the computational cost of these operations renders them unsuitable for token-level routing, where efficiency is critical.
In contrast, our method introduces compressed experts, which directly augment hidden state activations. This eliminates the need for dynamic parameter merging. Our approach retains compatibility with existing MoE architectures while offering greater flexibility, as compressed experts avoid the latency penalties of prior merging strategies.
Parameter Efficient Finetuning (PEFT)
@@
With the increasing sizes of LLMs, PEFT (He et al., 2021) has gained attention due to its capabilities to efficiently tune the model for specific downstream tasks. Instead of tuning the full model, PEFT methods introduce lightweight modules to augment the model, and only update those modules. Popular PEFT approaches include adapters (Houlsby et al., 2019), prompt tuning (Lester et al., 2021), LoRA (Hu et al., 2022) and (IA)3(Liu et al., 2022).
@@
Despite the similarities, it is important to note that compressed expert is not a PEFT method. While both approaches leverage lightweight modules, compressed experts focus on reducing computational costs during inference. The compressed experts serve as a compact representation of the auxiliary experts, while the main experts are used in their original form. Specifically, during SFT, the main experts are jointly trained with compressed experts. This design preserves the expressiveness of the main experts while reducing inference costs, addressing a fundamental efficiency-performance trade-off in MoE models.

### v2

@@ 1 Introduction
Figure 1: The inference time of MoE models grows linearly with increasing number of activated experts. The plot is generated using OLMoE architecture. Details on the inference speed measurement methodology are provided in Section˜3.1.
In practice, MoE models often require multiple activated experts to achieve desirable performance. For instance, Mixtral (Jiang et al., 2024) and Phi-MoE (Abdin et al., 2024) activate two experts, Qwen-MoE (Yang et al., 2024) activates four and OLMoE (Muennighoff et al., 2024) activates eight. The efficiency of MoE models largely depends on the number of activated experts. Specifically, the inference cost of MoE models increases linearly with this number (as shown in Figure˜1). Fewer activated experts reduce training and deployment costs as a smaller portion of the network is used. Conversely, activating more experts increases utilization of model capacity, often leading to superior performance. Thus, the selection of the number of activated experts poses a fundamental trade-off between efficiency and performance.
@@ 2.1 MoE Layers
where $\alpha_{i}=\text{top-}k(\text{Softmax}(R(h)))_{i}$. In this formulation, the top-$k$ mechanism ensures sparse activation, where only the $k$ experts with the highest routing weights contribute to the output, while the remaining nkn-k experts are effectively disabled ($\alpha_{i}=0$). Typically, only a small fraction of experts is utilized during each forward pass, i.e., $k\ll n$, enabling substantial computational savings. For instance, Phi-MoE (Abdin et al., 2024) activates 2 out of 16 experts, and OLMoE (Muennighoff et al., 2024) activates 8 out of 64 experts, etc. This sparse activation significantly reduces computational costs compared to dense models, making MoE layers efficient and scalable.
@@ 2.2 Compressed Experts
Conventionally, for each input token, $k$ out of the $n$ experts are activated. We can categorize these activated experts into $k_{m}$ main experts and $k_{a}=k-k_{m}$ auxiliary experts, based on their ordered routing weights. In many scenarios, instead of fully activating all $k$ experts, only forwarding the main experts can lead to a better performance-cost trade-off. However, we do not want to completely discard the information in auxiliary experts. Compressed experts provide a solution to leverage their compressed counterparts to reduce computational cost while retaining their contributions. We introduce it with the following three steps (illustrated in Figure˜3).
@@
Hidden states augmentation: The aggregated compressed expert $\theta$ is incorporated into the hidden state $h$ through an element-wise product. This operation has proven highly effective by PEFT methods such as (IA)3 (Liu et al., 2022), which applies a similar transformation to attention activations. Then, the augmented hidden state is passed only through the main experts, eliminating the need for forward computation through the $k_{a}$ auxiliary experts, thereby reducing inference cost. Note that using an alternative approach that modifies model parameters dynamically would introduce substantial complexity, as it would require constructing a new model configuration for each token in a sequence on the fly, given the combinatorial number of possible expert selections (i.e., choosing k active experts from n total experts). The element-wise product avoids this issue by enabling seamless integration of compressed experts without disrupting token-level independence.
@@ 2.3 Expert Reduction
While compressed experts are flexible and can be applied to any number of main and auxiliary experts, we investigate the conditions under which they are most effective. Specifically, we use OLMoE, which has 8 active experts ($k=8$), as a case study. We evaluate different configurations by varying the number of main experts ($k_{m}\in\{1,2,4,8\}$), and replace the remaining $k_{a}=k-k_{m}$ auxiliary experts with compressed experts (abbreviated as CE). To assess the impact of compressed experts in downstream adaptation, we conduct experiments using supervised finetuning (SFT) on pretrained OLMoE with mathematical data, followed by evaluation on GSM8k with 0-shot CoT (details provided in Section˜3.1).
@@
As shown in Table˜1, compressed experts consistently improve performance across reduced configurations, demonstrating their ability to recover performance lost due to expert reduction. However, their effectiveness is limited when more than half of the experts are treated as auxiliary.
@@ 2.4 Reduction in Activate Parameters
Replacing half of the activated experts with compressed experts significantly reduces the number of active parameters. Taking Phi-MoE as an example, each MoE layer consists of three weight matrices of size $4096\times 6400$. The total non-MoE parameters are approximately 2.4B. In the original configuration with 2 activated experts across 32 MoE layers, the total number of MoE parameters is $3\times 4096\times 6400\times 32\times 2\approx 5.03B.$ With compressed experts, only 1 expert is activated per layer, and the compressed experts add minimal parameters (underlined below): $3\times(4096\times 6400\times 1+\underline{2\times 4096+6400})\times 32\approx 2.52B$. In total, in one forward pass, the number of active parameters for the original configuration is 7.45B, while the one with compressed experts is only 4.93B, resulting in a saving of 33.8%. We provide a similar calculation for OLMoE in Appendix˜A, which results in a 31.4% reduction in active parameters.
Figure 4: The performance of Phi-MoE versus the inference latency, each point representing a different expert configuration. The Top-1 w/ CE configuration performs closely to the Top-2 configuration while achieving low inference latency close to Top-1. A similar plot for OLMoE is in Appendix˜B.
@@ 2.5 Performance-Latency Trade-off
To analyze the effectiveness of compressed experts, we compare the performance and inference latency of three configurations: top-2, top-1 with compressed experts (abbreviated as CE), and top-1, as shown in Figure˜4. The performance is evaluated on the Phi-MoE finetuned on mathematical reasoning data, with detailed training and evaluation procedures described in Section˜3.1.
@@
Introducing compressed experts in the top-1 w/ CE configuration effectively addresses this trade-off. With a latency of 4.35 seconds, the compressed expert configuration adds only a minimal computational overhead compared to top-1. However, it achieves a substantial performance improvement over top-1, closing much of the gap between top-1 and top-2 by reaching a performance of 57.4%. This demonstrates the ability of compressed experts to augment the utility of the top-1 expert with minimal additional cost, making it a practical and efficient model editing approach.
@@ 3 Experiments
In Section˜3.1, we detail the experimental setup, including model configurations, datasets and training specifics. In Section˜3.2, we first present the performance of compressed experts on evaluation benchmarks, then show their the inference cost savings, demonstrating that compressed experts achieves an effective balance between performance and efficiency. Finally, in Section˜3.3, we provide a detailed analysis on the task-specific performance of compressed experts.
@@ 3.1 Setup
Models. We evaluate our approach on two latest and representative MoE models: Phi-MoE (Abdin et al., 2024) and OLMoE (Muennighoff et al., 2024). These models vary in size and the number of activated experts, as summarized in Table˜2.
Task $(\rightarrow)$
Metric $(\rightarrow)$
@@
Task $(\rightarrow)$
Metric $(\rightarrow)$
@@
Task $(\rightarrow)$
Metric $(\rightarrow)$
@@
Task $(\rightarrow)$
Metric $(\rightarrow)$
@@
Training datasets. To evaluate compressed experts in various domains and applications, we conduct experiments on three tasks with different scopes. For specialized tasks including mathematics and coding, we finetune models on MathInstruct (Yue et al., 2024) and Magicoder (Wei et al., 2023). For general capabilities, we finetune models on TÜLU 3 (Lambert et al., 2024), a general-purpose instruction-following dataset targeting skills including reasoning, math, coding and safety.
Evaluation datasets. For evaluation, we benchmark models on both specialized and general tasks. For mathematical reasoning and coding, we evaluate on GSM8K (Cobbe et al., 2021) and HumanEval (Chen et al., 2021) respectively. To assess general-purpose capabilities, we further include IFEval (Zhou et al., 2023), TruthfulQA (Lin et al., 2021), BBH (Suzgun et al., 2023). Note that for HumanEval, we use Pass@1 for Phi-MoE and Pass@10 for OLMoE respectively. Since OLMoE is a smaller model with inherently lower capacity for complex coding tasks, Pass@10 provides finer-grained resolution to better capture performance differences across various configurations. The detailed evaluation metric for each task are presented in the tables shown below.
Baselines. Following the methodology outlined in Section˜2.5, we replace half of the activated experts with compressed experts and compare their performance against two baselines: the pretrained model and the configuration with half of the activated experts. Additionally, we report the performance of the full-expert configuration as a performance oracle for reference.
Inference latency evaluation. To evaluate the reduction of inference time with compressed experts, we follow the practice in the latency benchmark presented in Kwon et al. (2023). We create dummy prompts of batch size 8 and sequence length 32 with randomly generated tokens. Then, we pass the dummy prompts to the model and let it generate completions. The model first goes through 10 warmup iterations, and the reported latency is averaged over 30 iterations of completions.
@@ 3.2 Main Results
We present the performance of compressed experts on general tasks and specialized tasks respectively. Table˜3 and Table˜4 present the performance on a diverse set of general evaluation tasks, along with the inference latency for each configuration. Table˜5 and Table˜6 present the performance of Phi-MoE and OLMoE on specialized tasks, including mathematic and coding. Beyond raw performance, we additionally report normalized performance: $\text{perf}_{\text{CE}}/\text{perf}_{\text{full}}$. It quantifies how well compressed experts approximate the performance of the full-expert configuration, providing a direct measure of performance retention after expert reduction.
Performance comparison. The results demonstrate that compressed experts significantly narrow the performance gap with fully activated expert configurations. Across all models and tasks, compressed experts consistently recover more than 90% of the full expert performance, showcasing their ability to retain key information from auxiliary experts. Additionally, the incorporation of compressed experts consistently outperform the halved baseline. This shows that information contained in auxiliary experts still has a meaningful contribution to the model performance. This allows MoE models to maintain strong task performance with fewer active parameters. Similar to the findings in Huang et al. (2024), we observe that the optimal number of activated experts varies based on task complexity and broadness, and compressed experts exhibit task-dependent efficacy, which we detail in Section˜3.3.
Inference latency. Augmenting halved configurations with compressed experts (CE) incurs only marginal overhead. For Phi-MoE, top-1 with CE increases latency by 8.5% over top-1, while OLMoE’s top-4 with CE adds 9.7% latency compared to vanilla top-4. This overhead arises from lightweight operations such as element-wise multiplications, which is orders of magnitude cheaper than executing a full expert. Critically, CE-augmented configurations remain far more efficient than their original counterparts: Phi-MoE top-1 with CE is 22% faster than top-2, and OLMoE top-4 with CE is 18.4% faster than top-8. Combined with their strong empirical performance, we comprehensively validate that the compressed experts effectively balance efficiency and performance.
@@ 3.3 Empirical Analysis
We provide a more detailed analysis on per-task performance in the general evaluation setting (Tables˜3 and 4). Unlike specialized evaluation with a focus on specific domains, this setting covers factual correctness, logical reasoning, and precise instruction-following abilities. The diversity of tasks makes it challenging for MoE models to optimize for every task simultaneously, which in turn affects how well compressed experts perform. We observe that not all tasks benefit equally from compressed experts. Here, we mainly base our analysis on the performance improvement from the halved baseline.
@@
Overall, compressed experts provide an effective balance between efficiency and performance, consistently recovering over 90% of the full expert performance across both models. The performance recovery could vary across different task types, with structured reasoning tasks benefiting more than knowledge-intensive or multi-step reasoning tasks.
@@ 4 Related Works
Mixture-of-Experts (MoE). Mixture-of-Experts (MoE) models have emerged as a powerful approach to scaling up model capacity while maintaining computational efficiency. By activating only a subset of experts during inference, MoE models significantly reduce computational overhead compared to dense models. This advantage has led to a surge in open-source MoE implementations, each varying in the number of activated experts per layer. For instance, Mixtral (Jiang et al., 2024), JetMoE (Shen et al., 2024), OpenMoE (Xue et al., 2024) and Phi-MoE (Abdin et al., 2024) activate two experts, Qwen-MoE (Yang et al., 2024) activates four, DeepSeekMoE (Dai et al., 2024) activates six and OLMoE (Muennighoff et al., 2024) activates eight. However, recent work by Huang et al. (2024) demonstrates that fully utilizing all activated experts can introduce redundancy for certain tasks, as not all tasks require the full capacity of the top-$k$ experts. Inspired by this insight, we propose compressed experts, designed to compress auxiliary experts during model editing. Compressed experts are model-agnostic and can be seamlessly integrated into existing MoE architectures, including those mentioned above. By reducing the number of activated experts while maintaining performance, our approach enhances computational efficiency without compromising model quality.
@@
Soft merging of experts in MoE. Inspired by the surge of model merging techniques (Ilharco et al., 2023; Wortsman et al., 2022; Yadav et al., 2023; Yang et al., 2023; He et al., 2025), recent works propose to merge activated experts for more precise gradient computation and faster inference. However, these approaches suffer from a common limitation: the merging operations introduce significant computational overhead, making them impractical for token-based routing. Specifically, those approaches require merging experts on-the-fly based on the per-token routing weights. For instance, SMEAR (Muqeeth et al., 2024) and Lory (Zhong et al., 2024) merge experts via averaging the expert parameters based on the routing weights. While effective in sequence-based or semantic-based routing scenarios, the computational cost of these operations renders them unsuitable for token-level routing, where efficiency is critical. In addition, MC-SMoE (Li et al., 2023) provides a static merged expert configuration shared across all inputs. While this reduces computational costs, it sacrifices the token-level routing flexibility that makes MoE models powerful. In contrast, our method introduces compressed experts, which directly augment hidden state activations. This eliminates the need for dynamic parameter merging. Our approach retains compatibility with existing MoE architectures while offering greater flexibility, as compressed experts avoid the latency penalties of prior merging strategies.
Model pruning. We highlight two key advantages of our approach over pruning (Lu et al., 2024; He et al., 2024; Xie et al., 2024), which prunes the MoE model at an expert level. i) Higher efficiency: Pruning is a post-processing method after full training. In contrast, our method integrates compressed experts during training, allowing them to learn from both training data and auxiliary experts. Additionally, besides saving inference, the lightweight structure enables our method to achieve an approximate 20% reduction in training time as well by eliminating redundant expert computations, whereas post-processing methods like pruning require standard training plus an additional compression step. ii) Better knowledge preservation: Pruning removes experts deemed less important, which can lead to the loss of potentially valuable knowledge, even from experts that contribute less overall (Section˜2.3). In contrast, our goal is to preserve the information, but with smaller capacity.
@@
Parameter Efficient Finetuning (PEFT). With the increasing sizes of LLMs, PEFT (He et al., 2021) has gained attention due to its capabilities to efficiently tune the model for specific downstream tasks. Instead of tuning the full model, PEFT methods introduce lightweight modules to augment the model, and only update those modules. Popular PEFT approaches include adapters (Houlsby et al., 2019), prompt tuning (Lester et al., 2021), LoRA (Hu et al., 2022) and (IA)3(Liu et al., 2022).
@@
Despite the similarities, compressed expert is not a PEFT method. While both approaches leverage lightweight modules, compressed experts focus on reducing computational costs during inference. The compressed experts serve as a compact representation of the auxiliary experts, while the main experts are used in their original form. Specifically, during SFT, the main experts are jointly trained with compressed experts. This design preserves the expressiveness of the main experts while reducing inference costs, addressing a fundamental efficiency-performance trade-off in MoE models.